OpenStudy (anonymous):

X2-2x-13=0 in the indian quadratic equation

OpenStudy (anonymous):

Can you describe the indian quadratic equation

OpenStudy (anonymous):

Quadratics: solvers Practice! Answers archive Lessons Word Problems In Depth This Lesson (THE INDIAN METHOD OF SOLVING A QUADRATIC EQUATION) was created by by Theo(3348) : View Source, Show About Theo: The Indian Method of solving a quadratic equation is as follows: step 1: Move the constant term to the right side of the equation step 2: Multiply each term in the equation by four times the coefficient of the x^2 term step 3: Square the coefficient of the original x term and add it to both sides of the equation step 4: Take the square root of both sides step 5: Set the left side of the equation to the positive square root of the number on the right side and solve for x step 6: Set the left side of the equation to the negative square root of the number on the right side of the equation and solve for x

OpenStudy (merchandize):

Method 1 x² - 2x - 13 = 0 x² - 2x + 1 – 14 = 0 x² - 2x + 1 = 14 (x – 1)² = 14 x – 1 = ± √14 x = 1 + √14 or x = 1 – √14